OpenStudy (anonymous):

The derivative of the inverse function of \(x^5 + 3 x^3 + 7 x + 2\) at the point (13,1) is?

OpenStudy (anonymous):

y = x^5 + 3x^3 + 7x + 2 We want dx/dy at the point (1, 13) y' = 5x^4 + 9x^2 + 7 thus dx/dy = 1/(5x^4 + 9x^2 + 7) at the point x = 1, dx/dy = 1/21 = 1/21